# Descriptive Image Quality Assessment in the Wild

### v2

@@ Abstract
With the rapid advancement of Vision Language Models (VLMs), VLM-based Image Quality Assessment (IQA) seeks to describe image quality linguistically to align with human expression and capture the multifaceted nature of IQA tasks. However, current methods are still far from practical usage. First, prior works focus narrowly on specific sub-tasks or settings, which do not align with diverse real-world applications. Second, their performance is sub-optimal due to limitations in dataset coverage, scale, and quality. To overcome these challenges, we introduce Depicted image Quality Assessment in the Wild (DepictQA-Wild). Our method includes a multi-functional IQA task paradigm that encompasses both assessment and comparison tasks, brief and detailed responses, full-reference and non-reference scenarios. We introduce a ground-truth-informed dataset construction approach to enhance data quality, and scale up the dataset to 495K under the brief-detail joint framework. Consequently, we construct a comprehensive, large-scale, and high-quality dataset, named DQ-495K. We also retain image resolution during training to better handle resolution-related quality issues, and estimate a confidence score that is helpful to filter out low-quality responses. Experimental results demonstrate that DepictQA-Wild significantly outperforms traditional score-based methods, prior VLM-based IQA models, and proprietary GPT-4V in distortion identification, instant rating, and reasoning tasks. Our advantages are further confirmed by real-world applications including assessing the web-downloaded images and ranking model-processed images. Datasets and codes will be released in our project page. Corresponding Author. Project Page: https://depictqa.github.io/depictqa-wild/.
1 Introduction
@@
Image Quality Assessment (IQA) aims to measure and compare the quality of images, expecting to align with human perception. With the emergence of Vision Language Models (VLMs) [26, 41, 68], VLM-based IQA begins to attract more research interest [61, 62, 64, 65, 71]. These methods leverage VLMs to describe image quality using language, recognizing that language better mirrors human expression, and captures the multifaceted nature of IQA tasks [71]. However, existing methods are still far from VLM-based IQA in the wild, especially in their functionality and performance.
@@
Functionality. There are various application scenarios of IQA, but existing VLM-based IQA models only support a few of them. For example, one scenario involves assessing a single image downloaded from the web, while another requires comparing multiple images handled by different algorithms. Also, image restoration needs to assess an image against a reference, while image generation requests non-reference assessments. Therefore, an in-the-wild IQA model should be multi-functional to cater to such diverse scenarios. However, existing methods limit to a specific subset of these tasks, such as single-image assessment [62], multi-image comparison [64], or full-reference setting [71], etc. Hence, the limitations in functionality hinder prior methods to be an IQA method in the wild.
@@
Performance. Many IQA methods perform well on some specific datasets but may generalize poorly to other images with different contents or distortions. For instance, Co-Instruct [64] performs well on TID2013 [43] (85.0%), but drops significantly to 50.7% when testing on BAPPS [77]. A more comprehensive comparison on our newly created benchmark is given in Fig. 1, where it shows that previous works [62, 64] under-perform even within their defined tasks and settings. One potential cause for this is the limited scope of their training datasets. For example, the added distortion category in Q-Instruct [62] is limited; Co-Instruct [64] directly utilizes GPT-4V [41], which is not accurate in IQA tasks, to generate data; and the dataset scale in DepictQA [71] remains small. Additionally, these methods are constrained in their usage by resizing images to a fixed resolution [62, 64], while the image resolution is critical for quality assessment. Therefore, the dataset’s coverage, quality, and scale together with the training techniques limit the performance of previous methods.
@@
Figure 1: Performance comparison. Our model surpasses previous works including Q-Instruct [62], Co-Instruct [64], and the proprietary GPT-4V [41] across a broad range of tasks in both full-reference and non-reference settings. Traditional score-based IQA methods like LPIPS [77] and MUSIQ [20] have no language abilities, and thus can only be used in instant rating task.
@@
To overcome these challenges, we build an in-the-wild, multi-functional IQA model. We group multi-functional IQA tasks into two types, as shown in Fig. 2. (a) Single-image assessment aims to evaluate the quality of a single image by identifying distortions (e.g., “blur” in Fig. 2a top). It can also analyze the distortions’ impacts on contents (e.g., blur “affecting the definition of mountains and trees” in Fig. 2a bottom). (b) Paired-image comparison focuses on comparing the quality of two distorted images based on the clarity, colorfulness, and sharpness of presented contents. For example, in Fig. 2b, despite reduced contrast, “Image A maintains more scene integrity”, as “Image B’s serious noise level is more detrimental”. We omit multi-image comparison since it is an easy extension of a pairwise one [15]. Each task comprises a brief sub-task focusing on fundamental abilities, and a detailed sub-task fostering reasoning capacities. Furthermore, all tasks accommodate both full-reference and non-reference settings. These designs cater to diverse application scenarios.
@@
Under the multi-functional task paradigm, we construct a new large-scale dataset, DQ-495K, for comprehensive and accurate training and evaluation. First, for diverse distortion, we design and implement 35 types of distortions, each with 5 levels. Second, to enhance the label quality, we inform GPT-4V of the low-level ground truths (e.g., distortion information) to leverage its strong high-level perception and language abilities, while avoiding its relatively sub-optimal IQA capabilities. Third, to increase the dataset scale, we scale up the data amount to 495K under the brief-detail combined framework [71]. Moreover, our dataset is suitable for both full-reference and non-reference settings.
@@
With DQ-495K dataset, we then train a VLM model, named Depicted image Quality Assessment in the Wild (DepictQA-Wild). During training, the original image resolution is retained, leading to a better quality perception regarding resolution. Furthermore, we estimate the confidence of responses from key tokens, providing vital auxiliary information, especially for filtering low-quality responses.
@@
The performance of DepictQA-Wild is evaluated in Fig. 1 and Sec. 5. In brief tasks, our model surpasses general VLMs, IQA-specific VLMs, and score-based IQA methods by a large margin. For example, we achieve 95.9% in non-reference comparison on TID2013, remarkably surpassing Co-Instruct (85.0%) and GPT-4V (88.1%). In detailed tasks, our model also excels, e.g., recording 74.9% in non-reference comparison reasoning, compared to 48.1% for Co-Instruct and 60.3% for GPT-4V. At last, experiments on real-world applications including assessing web-downloaded images and comparing model-restored images further demonstrate our superiority. We hope that our multi-functional model could serve as a stepping stone towards a unified VLM-based IQA model. Although not yet fully realized, our method showcases the potential of VLM-based IQA models.
@@
Figure 2: Task paradigm and qualitative results. DepictQA-Wild focuses on two tasks including single-image assessment and paired-image comparison in both full-reference and non-reference settings. Each task contains a brief sub-task focusing on the fundamental IQA ability, and a detailed sub-task fostering the reasoning capacities. The shown results are tested in the non-reference setting.
2 Related Works
@@
Score-based IQA methods. Traditional IQA methods rely on scores to assess image quality and can be divided into full-reference and non-reference methods. (a) Full-reference methods compute a similarity score between a distorted image and a high-quality reference. Early works rely on human-designed metrics such as image information [49], structural similarity [59], phase congruency with gradient magnitude [75], etc. The rapid advancement of deep learning has also inspired learning-based IQA methods that measure image quality through data-driven training. Pioneered by PieAPP [44] and LPIPS [77], data-driven approaches [3, 4, 10, 11, 13, 69, 83] have spurred innovations in IQA, exhibiting high consistency with human judgments. (b) Non-reference methods directly regress a quality score without a reference image. Initially, human-designed natural image statistics are adopted [33, 36, 37, 38, 39, 48, 54]. Subsequently, deep-learning-based methods [19, 29, 42, 51, 52, 82, 85] replace hand-crafted statistics by learning quality priors from extensive data. Recent works focus on enhancing performance by introducing multi-scale features [20], CLIP pre-training [58], multi-dimension attention [66], continual learning [79], multitask learning [80], and so on. However, as discussed in [71], score-based IQA methods limit themselves in complex analyses and multi-aspect weighing of IQA, since the information provided by a single score is far from sufficient.
@@
Vision Language Models (VLMs) incorporate visual modality into large language models [8, 40, 55], aiming to leverage their emergent ability to achieve general visual ability. These VLMs [2, 9, 26, 41, 60, 67, 70, 76, 78, 84] have demonstrated a general visual ability and can tackle a variety of multi-modality tasks, including image captioning [1, 5, 72], visual question answering [14, 30, 31, 32], document understanding [34, 35, 50], etc. Although proficient in these high-level perception tasks, we demonstrate in Sec. 5 that general-purpose VLMs still struggle with IQA tasks.
@@
VLM-based IQA methods aim to achieve better alignment with human perception leveraging the power of VLMs [65]. Q-Bench [61, 81] establishes a comprehensive benchmark for evaluating general-purpose VLMs in low-level perception tasks. [86] evaluates various VLMs on the widely-adopted two-alternative forced choice (2AFC) task. Q-Instruct [62] enhances the low-level perception ability of VLMs by introducing a large-scale dataset. Q-Align [63] employs discrete text-defined levels for more accurate quality score regression. Co-Instruct [64] concentrates on the quality comparison among multiple images. DepictQA [71] performs quality description, quality comparison, and comparison reasoning in the full-reference setting. Nonetheless, as highlighted in Sec. 1, these methods focus only on specific aspects of IQA tasks, diverging from the original intents of VLMs’ universality and practical usage requirements, and their performance remains sub-optimal.
3 Task Paradigm and Dataset Construction
3.1 Task Paradigm
@@
As highlighted in the introduction, there are various application scenarios for IQA models. First, the evaluation objective can be either single-image assessment or paired-image comparison. The former is useful to rate a web-downloaded image, while the latter suits comparing images processed by two different algorithms. Second, the reference setting may be full-reference or non-reference. For example, image restoration requires assessments based on references, while image generation needs non-reference evaluations. Third, the response could be either brief or detailed. Brief responses suit well-targeted tasks (e.g., comparison without reasons), while detailed responses enhance interpretability and human interaction. To cater to such diverse scenarios, an in-the-wild IQA method should be multi-functional. Therefore, we aim to establish such a multi-functional task paradigm for VLM-based IQA research. As shown in Fig. 2, we focus on two tasks, each containing both brief and detailed sub-tasks, and supporting both full-reference and non-reference settings.
@@
  • Task2: paired-image comparison. (a) Brief sub-task: instant rating. Given two distorted images, the model should find the image with better quality. (b) Detailed sub-task: comparison reasoning. Building upon the comparison results, the model should first compare the content loss caused by distortions in the two images, then weigh different aspects to draw inferences, and finally justify its comparison results. Note that we omit the multi-image ($>$2) comparison since it can be achieved easily as the extension of paired case [15].
Compared with previous works, our design unifies various tasks, response types, and reference settings into a multi-functional paradigm. In contrast, Q-Instruct [62] focuses on non-reference single-image assessment, Co-Instruct [64] targets comparison among multiple images in the non-reference setting, and DepictQA [71] primarily addresses the full-reference setting. Although one can achieve unified IQA by combining these task-specific IQA models, it is impractical due to the significant increase in network parameters, considering that current VLMs are already quite large.
3.2 Distortion Library
Table 1: Overview of our distortion library with 12 super-categories and 35 sub-categories in total.
@@ 3.2 Distortion Library
Existing IQA datasets (e.g., BAPPS [77], PieAPP [44]) usually introduce distortions (e.g., noise, blur) into high-quality reference images to create distorted images for evaluation. However, these datasets do not publicly release the distortion information of each image, and their distortions only cover limited scenes. Therefore, we aim to develop a comprehensive large-scale distortion library.
@@
Distortion generation. Our distortion system comprises 12 super-categories in total, with each super-category consisting of multiple sub-categories. For instance, the “blur” category encompasses “Gaussian blur”, “motion blur”, “lens blur”, etc. In total, there are 35 sub-categories. For each sub-category, there are 5 severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. A summary is illustrated in Tab. 1. Considering the need to assess high-quality images as well, we retain the original image without any distortions in 5% proportion. See details in Appendix.
@@
Multi-distortion setups. In practical usage, multiple distortions may occur simultaneously on the same image. While a simple way to simulate them is to add multiple distortions recursively, real-world scenarios are more complex. First, one distortion may weaken another, such as “brighten” weakens “darken”, “blur” weakens “over-sharpen”. Second, certain distortions exhibit similar visual results, such as “pixelate” looks similar to “blur”, making it challenging to identify both if they are applied simultaneously. We also observe that humans can identify at most two distortions when three or more are applied. Hence, we limit the distortion number to two and manually review all possible combinations to exclude contradictory or similar combinations. See details in Appendix.
3.3 Dataset Construction

### v3

# Enhancing Descriptive Image Quality Assessment with A Large-scale Multi-modal Dataset

@@ Abstract
With the rapid advancement of Vision Language Models (VLMs), VLM-based Image Quality Assessment (IQA) seeks to describe image quality linguistically to align with human expression and capture the multifaceted nature of IQA tasks. However, current methods are still far from practical usage. First, prior works focus narrowly on specific sub-tasks or settings, which do not align with diverse real-world applications. Second, their performance is sub-optimal due to limitations in dataset coverage, scale, and quality. To overcome these challenges, we introduce the enhanced Depicted image Quality Assessment model (DepictQA-Wild). Our method includes a multi-functional IQA task paradigm that encompasses both assessment and comparison tasks, brief and detailed responses, full-reference and non-reference scenarios. We introduce a ground-truth-informed dataset construction approach to enhance data quality, and scale up the dataset to 495K under the brief-detail joint framework. Consequently, we construct a comprehensive, large-scale, and high-quality dataset, named DQ-495K. We also retain image resolution during training to better handle resolution-related quality issues, and estimate a confidence score that is helpful to filter out low-quality responses. Experimental results demonstrate that DepictQA-Wild significantly outperforms traditional score-based methods, prior VLM-based IQA models, and proprietary GPT-4V in distortion identification, instant rating, and reasoning tasks. Our advantages are further confirmed by real-world applications including assessing the web-downloaded images and ranking model-processed images. Codes, datasets, and model weights have been released in https://depictqa.github.io/.
I Introduction
@@
Image Quality Assessment (IQA) aims to measure and compare the quality of images, expecting to align with human perception. With the emergence of Vision Language Models (VLMs) [llava, gpt4v, mplug-owl2], VLM-based IQA begins to attract more research interest [qbench, qinstruct, coinstruct, iqasurvey, depictqa]. These methods leverage VLMs to describe image quality using language, recognizing that language better mirrors human expression, and captures the multifaceted nature of IQA tasks [depictqa]. However, existing VLM-based IQA methods still fall short especially in aspects of functionality and performance.
@@
Functionality. There are various application scenarios of IQA, but existing VLM-based IQA models only support a few of them. For example, one scenario involves assessing a single image downloaded from the web, while another requires comparing multiple images handled by different algorithms. Also, image restoration needs to assess an image against a reference, while image generation requests non-reference assessments. Therefore, a superior IQA model should be multi-functional to cater to such diverse scenarios. However, existing methods limit to a specific subset of these tasks, such as single-image assessment [qinstruct], multi-image comparison [coinstruct], or full-reference setting [depictqa], etc. Hence, the limitations in functionality hinder the wide applications of prior methods.
@@
Performance. Many IQA methods perform well on some specific datasets but may generalize poorly to other images with different contents or distortions. For instance, Co-Instruct [coinstruct] performs well on TID2013 dataset  [tid2013] (85.0%), but drops significantly to 50.7% when testing on BAPPS dataset [lpips]. A more comprehensive comparison on our newly created benchmark is given in Fig. 1, where it shows that previous works [qinstruct, coinstruct] under-perform even within their defined tasks and settings. One potential cause for this is the limited scope of their training datasets. For example, the added distortion category in Q-Instruct [qinstruct] is limited; Co-Instruct [coinstruct] directly utilizes GPT-4V [gpt4v], which is not accurate in IQA tasks, to generate data; and the dataset scale in DepictQA [depictqa] remains small. Besides, these methods are constrained in their usage by resizing images to a fixed resolution [qinstruct, coinstruct], while the image resolution is critical for quality. Therefore, the dataset’s coverage, quality, and scale together with the training techniques limit the performance of previous methods.
@@
Figure 1: Performance comparison. Our model surpasses previous works including Q-Instruct [qinstruct], Co-Instruct [coinstruct], and the proprietary GPT-4V [gpt4v] across a broad range of tasks in both full-reference and non-reference settings. Traditional score-based IQA methods like LPIPS [lpips] and MUSIQ [musiq] have no language abilities, and thus can only be used in instant rating task. Q-Instruct is only tested on single-image input tasks.
@@
To address these challenges, we propose a multi-functional IQA model to handle various image quality assessment tasks. We categorize these tasks into two types, as shown in Fig. 2. (a) Single-image assessment aims to evaluate the quality of a single image by identifying distortions (e.g., “blur” in Fig. 2a top). It can also analyze the distortions’ impacts on contents (e.g., blur “affecting the definition of mountains and trees” in Fig. 2a bottom). (b) Paired-image comparison focuses on comparing the quality of two distorted images based on the clarity, colorfulness, and sharpness of presented contents. For example, in Fig. 2b, despite reduced contrast, “Image A maintains more scene integrity”, as “Image B’s serious noise level is more detrimental”. We omit multi-image comparison since it is an easy extension of a pairwise one [pipal]. Each type includes basic brief sub-tasks for fundamental assessments and detailed sub-tasks to enhance reasoning abilities. Moreover, the model supports both full-reference and non-reference settings, making it adaptable to diverse scenarios.
@@
Under the multi-functional task paradigm, we construct a new large-scale dataset, DQ-495K, for comprehensive and accurate training and evaluation. First, for diverse distortion, we design and implement 35 types of distortions, each with 5 levels. Second, to enhance the label quality, we inform GPT-4V of the low-level ground truths (e.g., distortions) to leverage its strong high-level perception and language abilities, while avoiding its sub-optimal IQA capabilities. Third, to increase the dataset scale, we scale up the data amount to 495K under the brief-detail combined framework [depictqa]. Also, our dataset is suitable for both full-reference and non-reference settings.
@@
With the above collected DQ-495K dataset, we then train an enhanced Depicted image Quality Assessment VLM model (DepictQA-Wild). During training, the original image resolution is retained, leading to a better quality perception regarding resolution. Furthermore, we estimate the confidence of responses from key tokens, providing vital auxiliary information, especially for filtering low-quality responses.
@@
The performance of DepictQA-Wild is evaluated in Fig. 1 and Sec. V. In brief tasks, our model surpasses general VLMs, IQA-specific VLMs, and score-based IQA methods by a large margin. For example, we achieve 95.9% in non-reference comparison on TID2013 dataset, remarkably surpassing Co-Instruct (85.0%) and GPT-4V (88.1%). In detailed tasks, our model also excels, e.g., recording 74.9% in non-reference comparison reasoning, compared to 48.1% for Co-Instruct and 60.3% for GPT-4V. At last, experiments on real-world applications including assessing web-downloaded images and comparing model-restored images further demonstrate our superiority. We hope that our multi-functional model could serve as a stepping stone towards a unified VLM-based IQA model. Although not yet fully realized, our method showcases the potential of VLM-based IQA models.
@@
Figure 2: Illustration of our task paradigm and qualitative results. Our DepictQA-Wild focuses on two main tasks including single-image assessment and paired-image comparison in both full-reference and non-reference settings. Each task contains a brief sub-task focusing on the fundamental IQA ability, and a detailed sub-task fostering the reasoning capacities. More qualitative results are provided in Fig. S5 and S7 of Supp. Mat.
II Related Works
@@
Score-based IQA methods. Traditional IQA methods rely on scores to assess image quality and can be divided into full-reference and non-reference methods. (a) Full-reference methods compute a similarity score between a distorted image and a high-quality reference. Early works rely on human-designed metrics such as image information [vif], structural similarity [ssim], phase congruency with gradient magnitude [fsim], etc. The rapid advancement of deep learning has also inspired learning-based IQA methods that measure image quality through data-driven training. Pioneered by PieAPP [pieapp] and LPIPS [lpips], data-driven approaches [WaDIQaM, JSPL, dists, A-DISTS, ghildyal2022stlpips, CVRKD, SRIF] have spurred innovations in IQA, exhibiting high consistency with human judgments. (b) Non-reference methods directly regress a quality score without a reference image. Initially, human-designed natural image statistics are adopted [ma2017learning, BRISQUE, niqe, moorthy2010two, DIIVINE, saad2012blind, tang2011learning]. Subsequently, deep-learning-based methods [CNNIQA, RankIQA, BPSQM, HyperIQA, graphiqa, CKDN, MetaIQA] replace hand-crafted statistics by learning quality priors from extensive data. Recent works focus on enhancing performance by introducing multi-scale features [musiq], CLIP pre-training [clipiqa], multi-dimension attention [maniqa], continual learning [zhang2022continual], multitask learning [liqe], and so on. However, as discussed in [depictqa], score-based IQA methods limit themselves in complex analyses and multi-aspect weighing of IQA, since the information provided by a single score is far from sufficient.
@@
Vision Language Models (VLMs) incorporate visual modality into Large Language Models (LLMs) [vicuna, gpt4, llama], aiming to leverage their emergent ability to achieve general visual ability. These VLMs [flamingo, instructblip, llava, gpt4v, vary, mplug-owl, lamm, zhang2023internlm, llama_adapter, minigpt4] have demonstrated a general visual ability and can tackle a variety of multi-modality tasks, including image captioning [nocaps, cococap, flickr], visual question answering [vqav2, mmbench, scienceqa], document understanding [chartqa, docvqa, textvqa], etc. Although proficient in these high-level visual perception tasks, we demonstrate in Sec. V that general-purpose VLMs still struggle with IQA tasks.
@@
VLM-based IQA methods aim to achieve better alignment with human perception leveraging the power of VLMs [iqasurvey]. Q-Bench [qbench] establishes a comprehensive benchmark for evaluating general-purpose VLMs in low-level perception tasks. [2afcprompt] evaluates various VLMs on the widely-adopted two-alternative forced choice (2AFC) task. Q-Instruct [qinstruct] enhances the low-level perception ability of VLMs by introducing a large-scale dataset. Q-Align [qalign] and DeQA-Score [deqascore] employ discrete text-defined levels for more accurate quality score regression. Co-Instruct [coinstruct] concentrates on the quality comparison among multiple images. DepictQA [depictqa] performs quality description, quality comparison, and comparison reasoning in the full-reference setting. Recently, inspired by the success of DeepSeek-R1 [deepseekr1], Q-Insight [qinsight] first introduces reasoning-induced quality regression by employing the GRPO reinforcement learning method [grpo] for IQA. Building upon this idea, VisualQuality-R1 [visualqualityr1] further improves IQA performance by incorporating pairwise ranking relationships into the reinforcement learning objective. Moreover, Q-Ponder [qponder] proposes a unified two-stage training framework, consisting of a cold-start pretraining phase followed by reinforcement learning-based fine-tuning. Nonetheless, as highlighted in Sec. I, most of these methods focus on specific aspects of IQA tasks, diverging from the original intents of VLMs’ universality and practical usage requirements.
Explainable AI for computer vision. Our descriptive IQA model could be seen as an explainable IQA approach, as it provides a human-understandable reasoning process. Recent studies in computer vision have increasingly emphasized explainability to enhance transparency and trust in deep models [xai_survey1, xai_transformer1, xai_transformer2, xai_cnn, xai_science]. Several efforts have also been made to design explainable IQA models [xai_iqa1, xai_iqa2, xai_iqa3]. However, most existing explainable IQA methods primarily rely on visualization-based interpretability (i.e., saliency or attribution maps) to highlight influential regions, which can qualitatively indicate what the model attends to but fails to fully capture the underlying reasoning or perceptual logic behind the score assignment. Recently, TIFA [xai_iqa_qa] leveraged language-based question answering to evaluate text-to-image results, demonstrating that linguistic explanations can serve as an intuitive bridge between human judgment and model understanding. Inspired by this, we take a further step to construct a descriptive IQA model that provides natural-language rationales to describe perceptual factors influencing the assessment.
III Task Paradigm and Dataset Construction
III-A Task Paradigm
@@
As highlighted in the introduction, there are various application scenarios for IQA models. First, the evaluation objective can be either single-image assessment or paired-image comparison. The former is useful to rate a web-downloaded image, while the latter suits comparing images processed by two different algorithms. Second, the reference setting may be full-reference or non-reference. For example, image restoration requires assessments based on references, while image generation needs non-reference evaluations. Third, the response could be either brief or detailed. Brief responses suit well-targeted tasks (e.g., comparison without reasons), while detailed responses enhance interpretability and human interaction. To cater to such diverse scenarios, a practical IQA method should be multi-functional. Therefore, we aim to establish such a multi-functional task paradigm for VLM-based IQA research. As shown in Fig. 2, we focus on two tasks, each containing both brief and detailed sub-tasks, and supporting both full-reference and non-reference settings.
@@
  • Task2: paired-image comparison. (a) Brief sub-task: instant rating. Given two distorted images, the model should find the image with better quality. (b) Detailed sub-task: comparison reasoning. Building upon the comparison results, the model should first compare the content loss caused by distortions in the two images, then weigh different aspects to draw inferences, and finally justify its comparison results. Note that we omit the multi-image (>>2) comparison since it can be achieved easily as the extension of paired case [pipal].
Compared with previous works, our design unifies various tasks, response types, and reference settings into a multi-functional paradigm. In contrast, Q-Instruct [qinstruct] focuses on non-reference single-image assessment, Co-Instruct [coinstruct] targets comparison among multiple images in the non-reference setting, and DepictQA [depictqa] primarily addresses the full-reference setting. Although one can achieve unified IQA by combining these task-specific IQA models, it is impractical due to the significant increase in network parameters, considering that current VLMs are already quite large.
III-B Distortion Library
TABLE I: Overview of our distortion library with 12 super-categories and 35 sub-categories in total.
@@ III-B Distortion Library
Existing IQA datasets (e.g., BAPPS [lpips], PieAPP [pieapp]) usually introduce distortions (e.g., noise, blur) into high-quality reference images to create distorted images for evaluation. However, these datasets do not publicly release the distortion information of each image, and their distortions only cover limited scenes. Therefore, we aim to develop a comprehensive large-scale distortion library.
@@
Distortion generation. Our distortion system comprises 12 super-categories in total, with each super-category consisting of multiple sub-categories. For instance, the “blur” category encompasses “Gaussian blur”, “motion blur”, “lens blur”, etc. In total, there are 35 sub-categories. For each sub-category, there are 5 severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. A summary is illustrated in Tab. I. Considering the need to assess high-quality images as well, we retain the original image without any distortions in 5% proportion. See details in Supp. Mat.
@@
Multi-distortion setups. In practical usage, multiple distortions may occur simultaneously on the same image. While a simple way to simulate them is to add multiple distortions recursively, real-world scenarios are more complex. First, one distortion may weaken another, such as “brighten” weakens “darken”, “blur” weakens “over-sharpen”. Second, certain distortions exhibit similar visual results, such as “pixelate” looks similar to “blur”, making it challenging to identify both if they are applied simultaneously. We also observe that humans can identify at most two distortions when three or more are applied, as illustrated in Fig. S2 of Supp. Mat. Hence, we limit the distortion number to two and manually review all possible combinations to exclude contradictory or similar combinations. See details of multi-distortion setting in Tab. S1 and Sec. -A of Supp. Mat.
III-C Dataset Construction